# Diffusion Policy Policy Optimization

### v1

@@ 1 Introduction
Figure 1: We introduce DPPO, Diffusion Policy Policy Optimization, that fine-tunes pre-trained Diffusion Policy using policy gradient. DPPO affords structured exploration and training stability during policy fine-tuning, and the fine-tuned policy exhibits strong robustness and generalization. DPPO improves policy performance across benchmarks, including ones with pixel observations and with long-horizon rollouts that have been very challenging to solve using previous RL methods.
Contribution 2 (Demonstration of DPPO’s Performance). We show that for fine-tuning a pre-trained Diffusion Policy, DPPO yields consistent and marked improvements in training stability and final policy performance in comparison to a range of alternatives, including those based on off-policy Q-learning [84, 27, 86, 59] and weighted regression [53, 55, 37], as well as common policy parameterizations such as Gaussian and Gaussian Mixture models.
@@
Contribution 4 (Tackling challenging robotic tasks and settings). We demonstrate the effectiveness of DPPO in different challenging robotic and control settings, including ones with pixel observations and long-horizon manipulation tasks with sparse reward. We also deploy a policy trained in simulation via DPPO on real hardware in zero-shot, which exhibits a remarkably small sim-to-real gap compared to the baseline.
@@ 2 Related Work
Policy optimization methods update an explicit representation of an RL policy — typically parameterized by a neural network — by taking gradients through action likelihoods. Following the seminal policy gradient (PG) method [85, 76], there have been a range of algorithms that further improve training stability and sample efficiency such as DDPG [43] and PPO [69]. PG methods have been broadly effective in training robot policies [3, 34, 38, 13], largely due to their training stability with high-dimensional continuous action spaces, as well as their favorable scaling with parallelized simulated environments. Given the challenges of from-scratch exploration in long horizon tasks, PG has seen great success in fine-tuning a baseline policy trained from demonstrations [61, 78, 54].
@@
As demonstration data are often limited and/or suboptimal, there have been many approaches proposed to improve the performance of diffusion-based policies. One popular approach has been to guide the diffusion denoising process using objectives such as reward signal or goal conditioning [35, 2, 42, 80, 11]. More recent work has explored the possibilities of techniques including Q-learning and weighted regression, either from purely offline estimation [12, 84, 17], and/or with online interaction [37, 27, 59, 86].
@@
RL techniques have been used to fine-tune diffusion models such as ones for text-to-image generation [19, 20, 6, 16, 81]. Clark et al. [16] treat the denoising process as an MDP, to which they apply policy gradient updates, and Black et al. [6] apply the PPO update to this formalism. We build upon these earlier findings by embedding the denoising MDP into the environmental MDP of the dynamics in control tasks, forming a two-layer “Diffusion Policy MDP”. Though Psenka et al. [59] have already shown how PG can be taken through Diffusion Policy by propagating PG through both MDPs, they conjecture that it is likely to be ineffective due to large action variance caused by the increased effective horizon induced from the denoising steps. Our results contravene this supposition for diffusion-based policies in the fine-tuning setting.
@@ 3 Preliminaries
We consider Markov Decision Process (MDP) [^1] More generally, we can view our environment as a Partially Observed Markov Decision Process (POMDP) where the agent’s actions depend on observations $o$ of the states $s$ (e.g., action from pixels). Our implementation applies in this setting, but we omit additional observations from the formalism to avoid notional clutter. $\mathcal{M}_{\textsc{Env}}:=(\mathcal{S},\mathcal{A},P_{0},P,R)$ with states $s\in\mathcal{S}$, actions $a\in\mathcal{A}$, initial state distribution $P_{0}$, transition probabilities $P$, and reward $R$.
At each timestep $t$, the agent (e.g., robot) observes the state $s_{t}\in\mathcal{S}$, takes an action $a_{t}\sim\pi(a_{t}\mid s_{t})\in\mathcal{A}$, transitions to the next state $s_{t+1}$ according to $s_{t+1}\sim P(s_{t+1}\mid s_{t},a_{t})$ while receiving the reward $R(s_{t},a_{t})$ [^2] For simplicity, we overload $R(\cdot,\cdot)$ to denote both the random variable reward and its distribution.. Fixing the MDP $\mathcal{M}_{\textsc{Env}}$, we let $\mathbb{E}^{\pi}$ (resp. $\operatorname{\mathbb{P}}^{\pi}$) denote the expectation (resp. probability distribution) over trajectories, $\tau=(s_{0},a_{0},...,s_{T},a_{T})$ with length $T+1$, with initial state distribution $s_{0}\sim P_{0}$ and transition operator $P$. We aim to train a policy to optimize the cumulative reward, discounted by a function $\gamma(\cdot)$, such that the agent receives:
@@
Diffusion Policy (DP; see [15]) is a policy $\pi_{\theta}$ parameterized by a DDPM which takes in $s$ as a conditioning argument, and parameterizes $p_{\theta}(a^{k-1}\mid a^{k},s)$ as in (3.3). DPs can be trained via behavior cloning by fitting the conditional noise prediction $\varepsilon_{\theta}(a^{k},s,k)$ to predict added noise. Notice that unlike more standard policy parameterizations such as unimodal Gaussian policies, DPs do not maintain an explicit likelihood of $p_{\theta}(a^{0}\mid s)$. In this work, we adopt the common practice of training DPs to predict an action chunk — a sequence of actions a few time steps (denoted $T_{a}$) into the future — to promote temporal consistency. For fair comparison, our non-diffusion baselines use the same chunk size.
@@ 4 DPPO: Diffusion Policy Policy Optimization
As noted in Section 1 and Section 2, there has been much attention devoted to fine-tuning Diffusion Policy for improved performance [86, 59], focusing primarily on off-policy Q-learning [84, 27] and/or weighted regression. DPPO takes a different approach: differentiate through action likelihood and apply a policy gradient (PG) update like (3.2).
@@ 4.1 A Two-Layer “Diffusion Policy MDP”
Figure 3: We treat the denoising process in Diffusion Policy as a MDP, and the whole environment episode can be considered as a chain of such MDPs. Now the entire chain (“Diffusion Policy MDP”, $\mathcal{M}_{\textsc{DP}}$) involves a Gaussian likelihood at each (denoising) step and thus can be optimized with policy gradient. Blue circle denotes the state and red circle denotes the action in $\mathcal{M}_{\textsc{DP}}$.
As observed in [16, 6] and [59], a denoising process can be represented as a multi-step MDP in which policy likelihoods can be obtained directly. We extend this formalism by embedding the Diffusion MDP into the environmental MDP, obtaining a larger “Diffusion Policy MDP” denoted $\mathcal{M}_{\textsc{DP}}$, visualized in Fig. 3. Below, we use the notation $\updelta$ to denote a Dirac distribution and $\otimes$ to denote a product distribution.
Recall the environment MDP $\mathcal{M}_{\textsc{Env}}:=(\mathcal{S},\mathcal{A},P_{0},P,R)$ in Section 3. The Diffusion MDP $\mathcal{M}_{\textsc{DP}}$ uses indices $\bar{t}(t,k)=tK+(K-k-1)$ corresponding to $(t,k)$, which increases in $t$ but (to keep the indexing conventions of diffusion) decreases lexicographically with $K-1\geq k\geq 0$. We write states, actions and rewards as,
@@
That is, the transition moves the denoised action $a_{t}^{k}$ at step $\bar{t}(t,k)$ into the next state when $k>0$, or otherwise progresses the environment MDP dynamics with $k=0$. The pure noise $a_{t}^{K}$ is considered part of the environment when transitioning at $k=0$. In light of (3.3), the policy in $\mathcal{M}_{\textsc{DP}}$ takes the form
@@
Fortunately, (4.3) is a Gaussian likelihood, which can be evaluated analytically and is amenable to the policy gradient updates (see also [59] for an alternative derivation):
@@ 4.2 Instantiating DPPO with Proximal Policy Optimization
We apply Proximal Policy Optimization (PPO) [70, 18, 32, 1], a popular improvement of the vanilla policy gradient update.
Definition 4.1 (Generalized PPO).
Consider a general MDP. Given an advantage estimator $\hat{A}(s,a)$, the PPO update is given by [70] the sample approximation to
@@
where $\varepsilon$, the clipping ratio, controls the maximum magnitude of policy change from the previous policy.
@@
since $\bar{R}(\bar{t})=0$ at $k>0$. This fact also obviates the need of estimating the value at $k>1$ and allows us to use the following denoising-discounted advantage estimator [^3] In practice, we use Generalized Advantage Estimation (GAE) [69] that better balances variance and bias in estimating the advantage. We present the simpler form here.:
@@
The denoising-discounting has the effect of downweighting the contribution of noisier steps (larger $k$) to the policy gradient (see study in Section B.1). Lastly, we choose the value estimator to only depend on the “$s$” component of $\bar{s}$:
@@
which we find leads to more efficient and stable training compared to also estimating the value of applying the denoised action $a_{t}^{k=1}$ (part of $\bar{s}_{\bar{t}(t,0)}$) as shown in Section B.1.
@@ 4.3 Best Practices for DPPO
Diffusion Policy often uses up to $K=100$ denoising steps with DDPM to better capture the complex data distribution of expert demonstrations. With DPPO, we can choose to fine-tune only a subset of the denoising steps instead, e.g., the last $K^{\prime}$ steps. Experimental results in Section B.1 shows this speeds up DPPO training and reduces GPU memory usage without sacrificing the final performance. Instead of fine-tuning the pre-trained model weights $\theta$, we make a copy $\theta_{\textsc{ft}}$$\theta$ is frozen and used for the early denoising steps, while $\theta_{\textsc{ft}}$ is used for the last $K^{\prime}$ steps and updated with DPPO.
@@
Although DDIM is typically used as a deterministic sampler by setting $\eta=0$ in (4.9), we can use $\eta>0$ for fine-tuning in order to provide exploration noise and avoid calculating a Gaussian likelihood with a Dirac distribution. In practice, we set $\eta=1$ for training (equivalent to applying DDPM [72]) and then $\eta=0$ for evaluation. We use DDIM sampling for our pixel-based experiments and long-horizon furniture assembly tasks, where the efficiency improvements are much desired.
@@
We use the cosine schedule for $\sigma_{k}$ introduced in [49], which was originally annealed to a small value on the order of $1\text{e}{-4}$ at $k=0$. In DPPO, the values of $\sigma_{k}$ also translate to the exploration noise that is crucial to training efficiency. Empirically, we find that clipping $\sigma_{k}$ to a higher minimum value (denoted $\sigma^{\text{exp}}_{\text{min}}$, e.g., $0.01-0.1$) when sampling actions helps exploration (see study in Section B.1). Additionally we clip $\sigma_{k}$ to be at least $0.1$ (denoted $\sigma^{\text{prob}}_{\text{min}}$) when evaluating the Gaussian likelihood $\log\bar{\pi}_{\theta}(\bar{a}_{\bar{t}}|\bar{s}_{\bar{t}})$, which improves training stability by avoiding large magnitude.
@@
We study both Multi-layer Perceptron (MLP) and UNet [66] as the policy heads in Diffusion Policy. MLP offers a simpler setup, and we find that it generally fine-tunes more stably with DPPO. Meanwhile, UNet, only applying convolutions to $a_{t}^{k}$, has the benefit of allowing pre-training and fine-tuning with different action chunk size $T_{a}$, e.g., $16$ and $8$. We find that DPPO benefits from pre-training with larger $T_{a}$ (better prediction) and fine-tuning with smaller $T_{a}$ (more amenable to policy gradient)  [^4] With fully-connected layers in MLP, empirically we find that using different chunk sizes for pre-training and fine-tuning with MLP leads to training instability..
@@ 5 Performance Evaluation of DPPO
We begin our investigation of DPPO by studying its performance in various popular RL and robotics benchmarking environments. We compare to other RL approaches for fine-tuning a Diffusion Policy (Section 5.1), next to other policy parameterizations (Section 5.2), and then in multi-stage manipulation tasks including hardware evaluation (Section 5.3), and conclude with ablations (Section 5.4). While our evaluations focus primarily on fine-tuning, we also present training-from-scratch results in Appendix B. Wall-clock times are reported and discussed in Appendix C; they are roughly comparable (and often faster) than other diffusion-based RL baselines, though can be up to 2×2$\times$2 × slower than other policy parameterization baselines. Full choices of training hyperparameters and additional training details are presented in Appendix D.
@@
We first consider three OpenAI Gym locomotion benchmarks [8] common in the RL literature: {Hopper-v2, Walker2D-v2, HalfCheetah-v2}. All policies are pre-trained with the medium-level datasets from D4RL [24] with state input and action chunk size $T_{a}=4$. We use the original dense reward setup in fine-tuning.
@@
Next we consider four simulated robot manipulation tasks from the Robomimic benchmark [48], {Lift, Can, Square, Transport}, ordered in increasing difficulty. These tasks are more representative of real-world robotic tasks, and Square and Transport (Fig. 4) are considered very challenging for RL training. Both state and pixel policy input are considered. State-based and pixel-based policies are pre-trained with 300 and 100 Multi-Human demonstrations provided by Robomimic, respectively. We consider $T_{a}=4$ for Can, Lift, and Square, and $T_{a}=8$ for Transport. They are then fine-tuned with sparse reward equal to $1$ upon task completion.
@@
Finally, we demonstrate solving longer-horizon, multi-stage robot manipulation tasks from the Furniture-Bench [28] benchmark in both simulation and reality. We evaluate the methods on three simulated furniture assembly tasks, {One-leg, Lamp, Round-table}, shown in Fig. 4 and described in detail in Section D.7. We consider two levels of randomness over initial state distribution, Low and Med, defined by the benchmark. State policy input is considered. All policies are pre-trained with 50 human demonstrations collected in simulation and $T_{a}=8$. They are then fine-tuned with sparse (indicator of task stage completion) reward. We also evaluate the zero-shot sim-to-real performance with One-leg.

### v3

@@ 1 Introduction
Figure 1: We introduce DPPO, Diffusion Policy Policy Optimization, that fine-tunes pre-trained Diffusion Policy using policy gradient. DPPO treats the denoising process as a Markov Decision Process (“Diffusion MDP”) allowing the task reward signal (from “Environment MDP“) to easily propagate through Diffusion Policy (Fig. 3). Policy gradient update then leverages the tractable Gaussian likelihood at each denoising step. Extensive experiments in simulation and hardware show DPPO affords structured exploration and training stability during policy fine-tuning, and the final policy exhibits strong robustness and generalization.
Contribution 2 (Demonstration of DPPO’s Performance). We show that for fine-tuning a pre-trained Diffusion Policy, DPPO yields strong training stability across tasks and marked improvements in final performance in challenging robotic tasks in comparison to a range of alternatives, including those based on off-policy Q-learning [100, 35, 103, 74] and weighted regression [67, 70, 47], other demo-augmented RL methods [7, 62, 40], as well as common policy parameterizations such as Gaussian and Gaussian Mixture models with PG updates.
@@
Contribution 4 (Tackling challenging robotic tasks and settings). We show DPPO is effective in various challenging robotic and control settings, including ones with pixel observations and long-horizon manipulation tasks with sparse reward. We also deploy a policy trained in simulation via DPPO on real hardware in zero-shot, which exhibits a remarkably small sim-to-real gap compared to the baseline.
@@ 2 Related Work
Policy optimization methods update an explicit representation of an RL policy — typically parameterized by a neural network — by taking gradients through action likelihoods. Following the seminal policy gradient (PG) method [101, 91], there have been a range of algorithms that further improve training stability and sample efficiency such as DDPG [54] and PPO [84]. PG methods have been broadly effective in training robot policies [4, 43, 48, 17], largely due to their training stability with high-dimensional continuous action spaces, as well as their favorable scaling with parallelized simulated environments. Given the challenges of from-scratch exploration in long-horizon tasks, PG has seen great success in fine-tuning a baseline policy trained from demonstrations [76, 93, 68]. Our experiments find DPPO performing on-policy PG often achieves stronger final performance in manipulation tasks, especially ones with long horizon and high-dimensional action space, than off-policy Q-learning methods [7, 62, 40]. See Section A.1 for extended discussions on PG fine-tuning of robot policies and related methods.
@@
As demonstration data are often limited and/or suboptimal, there have been many approaches proposed to improve the performance of diffusion-based policies. One popular approach has been to guide the diffusion denoising process using objectives such as reward signal or goal conditioning [45, 2, 53, 96, 15]. More recent work has explored the possibilities of techniques including Q-learning and weighted regression, either from purely offline estimation [16, 100, 22], and/or with online interaction [47, 35, 74, 103]. See Section A.2 for detailed descriptions of these methods.
@@
RL techniques have been used to fine-tune diffusion models such as ones for text-to-image generation [24, 25, 9, 97]. Black et al. [9] treat the denoising process as an MDP and apply PPO updates. We build upon these earlier findings by embedding the denoising MDP into the environmental MDP of the dynamics in control tasks, forming a two-layer “Diffusion Policy MDP”. Though Psenka et al. [74] have already shown how PG can be taken through Diffusion Policy by propagating PG through both MDPs, they conjecture that it is likely to be ineffective due to large action variance caused by the increased effective horizon induced from the denoising steps. Our results contravene this supposition for diffusion-based policies in the fine-tuning setting.
@@ 3 Preliminaries
We consider a Markov Decision Process (MDP) [^1] More generally, we can view our environment as a Partially Observed Markov Decision Process (POMDP) where the agent’s actions depend on observations $o$ of the states $s$ (e.g., action from pixels). Our implementation applies in this setting, but we omit additional observations from the formalism to avoid notional clutter. $\mathcal{M}_{\textsc{Env}}:=(\mathcal{S},\mathcal{A},P_{0},P,R)$ with states $s\in\mathcal{S}$, actions $a\in\mathcal{A}$, initial state distribution $P_{0}$, transition probabilities $P$, and reward $R$.
At each timestep $t$, the agent (e.g., robot) observes the state $s_{t}\in\mathcal{S}$, takes an action $a_{t}\sim\pi(a_{t}\mid s_{t})\in\mathcal{A}$, transitions to the next state $s_{t+1}$ according to $s_{t+1}\sim P(s_{t+1}\mid s_{t},a_{t})$ while receiving the reward $R(s_{t},a_{t})$ [^2] For simplicity, we overload $R(\cdot,\cdot)$ to denote both the random variable reward and its distribution.. Fixing the MDP $\mathcal{M}_{\textsc{Env}}$, we let $\mathbb{E}^{\pi}$ (resp. $\operatorname{\mathbb{P}}^{\pi}$) denote the expectation (resp. probability distribution) over trajectories $(s_{0},a_{0},...,s_{T},a_{T})$ with length $T+1$, with initial state distribution $s_{0}\sim P_{0}$ and transition operator $P$. We aim to train a policy to optimize the cumulative reward, discounted by a function $\gamma(\cdot)$: :
@@
Diffusion Policy (DP; see Chi et al. [20]) is a policy $\pi_{\theta}$ parameterized by a DDPM which takes in $s$ as a conditioning argument, and parameterizes $p_{\theta}(a^{k-1}\mid a^{k},s)$ as in Eq. (3.3). DPs can be trained via behavior cloning by fitting the conditional noise prediction $\varepsilon_{\theta}(a^{k},s,k)$ to predict the added noise. Notice that unlike more standard policy parameterizations such as unimodal Gaussian policies, DPs do not maintain an explicit likelihood of $p_{\theta}(a^{0}\mid s)$. In this work, we adopt the common practice of training DPs to predict an action chunk — a sequence of actions a few time steps (denoted $T_{p}$) into the future — to promote temporal consistency. Policy executes $T_{a}\leq T_{p}$ steps of the predicted chunk before the next prediction. From now on $a\in\mathcal{A}$ denotes the entire executed action chunk. For fair comparison, our diffusion and non-diffusion baselines use the same $T_{a}$ unless noted.
@@ 4 DPPO: Diffusion Policy Policy Optimization
As noted in Section 1 and Section 2, there has been much attention devoted to fine-tuning Diffusion Policy for improved performance [103, 74], focusing primarily on off-policy Q-learning [100, 35] and/or weighted regression. DPPO takes a different approach: differentiate through action likelihood and apply a policy gradient (PG) update like Eq. (3.2).
@@ 4.1 A Two-Layer “Diffusion Policy MDP”
Figure 3: We treat the denoising process in Diffusion Policy as an MDP, and the whole environment episode can be considered as a chain of such MDPs. Now the entire chain (“Diffusion Policy MDP”, $\mathcal{M}_{\textsc{DP}}$) involves a Gaussian likelihood at each (denoising) step and thus can be optimized with policy gradient. Blue circle denotes the state and red circle denotes the action in $\mathcal{M}_{\textsc{DP}}$.
As observed in [9] and [74], a denoising process can be represented as a multi-step MDP in which policy likelihood of each denoising step can be obtained directly. We extend this formalism by embedding the Diffusion MDP into the environmental MDP, obtaining a larger “Diffusion Policy MDP” denoted $\mathcal{M}_{\textsc{DP}}$, visualized in Fig. 3. Below, we use the notation $\updelta$ to denote a Dirac distribution and $\otimes$ to denote a product distribution.
Recall the environment MDP $\mathcal{M}_{\textsc{Env}}:=(\mathcal{S},\mathcal{A},P_{0},P,R)$ in Section 3. The Diffusion MDP $\mathcal{M}_{\textsc{DP}}$ uses indices $\bar{t}(t,k)=tK+(K-k-1)$ corresponding to $(t,k)$, which increases in $t$ but (to keep the indexing conventions of diffusion) decreases lexicographically with $K-1\geq k\geq 0$. The states, actions and rewards are
@@
That is, the transition moves the denoised action $a_{t}^{k}$ at step $\bar{t}(t,k)$ into the next state when $k>0$, or otherwise progresses the environment MDP dynamics with $k=0$. The pure noise $a_{t}^{K}$ is considered part of the environment when transitioning at $k=0$. In light of Eq. (3.3), the policy in $\mathcal{M}_{\textsc{DP}}$ takes the form
@@
Fortunately, Eq. (4.3) is a Gaussian likelihood, which can be evaluated analytically and is amenable to the policy gradient updates333[74] proposes a similar derivation but does not consider the denoising process as a MDP. See further clarification in Appendix A.:
@@ 4.2 Instantiating DPPO with Proximal Policy Optimization
We apply Proximal Policy Optimization (PPO) [85, 23, 41, 1], a popular improvement of the vanilla policy gradient update. The full pseudocode with implementation details are shown in Algorithm 1, Appendix B.
Definition 4.1 (Generalized PPO, clipping variant).
Consider a general MDP. Given an advantage estimator $\hat{A}(s,a)$, the PPO update [85] is the sample approximation to
@@
where $\varepsilon$, the clipping ratio, controls the maximum magnitude of the policy updated from the previous policy.
@@
since $\bar{R}(\bar{t})=0$ at $k>0$. This fact also obviates the need of estimating the value at $k>1$ and allows us to use the following denoising-discounted advantage estimator [^4] In practice, we use Generalized Advantage Estimation (GAE, Schulman et al. [84]) that better balances variance and bias in estimating the advantage. We present the simpler form here.:
@@
The denoising-discounting has the effect of downweighting the contribution of noisier steps (larger $k$) to the policy gradient (see study in Section C.2). Lastly, we choose the value estimator to only depend on the “$s$” component of $\bar{s}$:
@@
which we find leads to more efficient and stable training compared to also estimating the value of applying the denoised action $a_{t}^{k=1}$ (part of $\bar{s}_{\bar{t}(t,0)}$) as shown in Section C.2.
@@ 4.3 Best Practices for DPPO
Diffusion Policy often uses up to $K=100$ denoising steps with DDPM to better capture the complex data distribution of expert demonstrations. With DPPO, we can choose to fine-tune only a subset of the denoising steps instead, e.g., the last $K^{\prime}$ steps. Experimental results in Section C.2 shows this speeds up DPPO training and reduces GPU memory usage without sacrificing the final performance. Instead of fine-tuning the pre-trained model weights $\theta$, we make a copy $\theta_{\textsc{ft}}$$\theta$ is frozen and used for the early denoising steps, while $\theta_{\textsc{ft}}$ is used for the last $K^{\prime}$ steps and updated with DPPO.
@@
Although DDIM is typically used as a deterministic sampler by setting $\eta=0$ in Eq. (4.9), we can use $\eta>0$ for fine-tuning in order to provide exploration noise and avoid calculating a Gaussian likelihood with a Dirac distribution. In practice, we set $\eta=1$ for training (equivalent to applying DDPM [87]) and then $\eta=0$ for evaluation. We use DDIM sampling for our pixel-based experiments and long-horizon furniture assembly tasks, where the efficiency improvements are much desired.
@@
We use the cosine schedule for $\sigma_{k}$ introduced in [63], which was originally annealed to a small value on the order of $1\text{e}{-4}$ at $k=0$. In DPPO, the values of $\sigma_{k}$ also translate to the exploration noise that is crucial to training efficiency. Empirically, we find that clipping $\sigma_{k}$ to a higher minimum value (denoted $\sigma^{\text{exp}}_{\text{min}}$, e.g., $0.01-0.1$) when sampling actions helps exploration (see study in Section C.2). Additionally we clip $\sigma_{k}$ to be at least $0.1$ (denoted $\sigma^{\text{prob}}_{\text{min}}$) when evaluating the Gaussian likelihood $\log\bar{\pi}_{\theta}(\bar{a}_{\bar{t}}|\bar{s}_{\bar{t}})$, which improves training stability by avoiding large magnitude.
@@
We study both Multi-layer Perceptron (MLP) and UNet [81] as the policy heads in Diffusion Policy. MLP offers a simpler setup, and we find that it generally fine-tunes more stably with DPPO. Meanwhile, UNet, only applying convolutions to $a_{t}^{k}$, has the benefit of allowing fine-tuning with $T_{a}<T_{p}$, e.g., $T_{p}=16$ and $T_{a}=8$. We find that DPPO benefits from pre-training with larger $T_{p}$ (better prediction) and fine-tuning with smaller $T_{a}$ (more amenable to policy gradient) [^5] With fully-connected layers in MLP, empirically we find that fine-tuning with $T_{a}<T_{p}$ leads to training instability..
@@ 5 Performance Evaluation of DPPO
We study the performance of DPPO in popular RL and robotics benchmarking environments. We compare to other RL methods for fine-tuning a Diffusion Policy (Section 5.1), to other demo-augmented RL methods (Section 5.2), to other policy parameterizations with PG updates (Section 5.3), and then in multi-stage manipulation tasks including hardware evaluation (Section 5.4), and conclude with ablations (Section 5.5). While our evaluations focus primarily on fine-tuning, we also present training-from-scratch results in Appendix C. Wall-clock times are reported and discussed in Appendix D; they are roughly comparable (often faster) than other diffusion-based RL baselines, much faster than other demo-augmented RL baselines, though can be up to 2×2$\times$2 × slower than other policy parameterizations. Full choices of training hyperparameters and additional training details are presented in Appendix E.
@@
We first consider three OpenAI Gym locomotion benchmarks [11] common in the RL literature: {Hopper-v2, Walker2D-v2, HalfCheetah-v2}. All policies are pre-trained with the full medium-level datasets from D4RL [29] with state input and action chunk size $T_{a}=4$. We use the original dense reward setup in fine-tuning.
Environments: Franka Kitchen.
We also consider the Franka-Kitchen environment first introduced in [32] involving a Franka arm completing four countertop tasks. We use the three datasets from D4RL [29]: {Kitchen-Complete-v0, Kitchen-Partial-v0, Kitchen-Mixed-v0}, containing demonstrations of varying levels of completeness. State input is considered and action chunk size $T_{a}=4$ is used for DPPO. We use the original sparse reward setup in fine-tuning.
@@
Next we consider four simulated robot manipulation tasks from the Robomimic benchmark [60], {Lift, Can, Square, Transport}, ordered in increasing difficulty. These tasks are more representative of real-world robotic tasks, and Square and Transport (Fig. 4) are considered very challenging for RL training. Both state and pixel policy input are considered. State-based and pixel-based policies are pre-trained with 300 and 100 demonstrations provided by Robomimic, respectively. We mostly consider the noisier Multi-Human (MH) data from Robomimic but also consider the cleaner Proficient-Human (PH) data in some experiments. We consider $T_{a}=4$ for Can, Lift, and Square, and $T_{a}=8$ for Transport. They are then fine-tuned with sparse reward equal to $1$ upon task completion.
@@
Finally, we demonstrate solving longer-horizon, multi-stage robot manipulation tasks from the Furniture-Bench [36] benchmark. We consider three simulated furniture assembly tasks, {One-leg, Lamp, Round-table}, shown in Fig. 4 and described in detail in Section E.8. We consider two levels of randomness over initial state distribution, Low and Med, defined by the benchmark. All policies are pre-trained with 50 human demonstrations collected in simulation and $T_{a}=8$. They are then fine-tuned with sparse (indicator of task stage completion) reward. We also evaluate the zero-shot sim-to-real performance with One-leg.